Suppose that the term structure today is given as follows: Time to maturity YTM 1 1.2%...

  1. Suppose that the term structure today is given as follows:

Time to maturity

YTM

1

1.2%

2

1.6%

3

2.1%

4

2.9%

Suppose your investment horizon is 4 years, and you are considering buying a coupon bond with par 1000, coupon rate 3%, paying once a year.

Assume you believe in the expectations hypothesis. What is your expected realized yield?
Answer #1

Bond price at par = 1000

coupon rate = 3% p.a , Time = 4years

Coupon amount = 1000 * 3% = 30

Present value of Bond @ YTM  = PV of interest + PV of maturity value of bond

= 30 / (1+0.012) + 30 / (1+0.012)(1+0.016) + 30 / (1+0.012)(1+0.016)(1+0.021) +

30 / (1+0.012)(1+0.016)(1+0.021)(1+0.029) + 1000 / (1+0.012)(1+0.016)(1+0.021)(1+0.029)

= 1040.90

Expected realized yield = (1040.90-1000) /1000

= 0.0409 or 4.09%